If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?

Answer by ewatrrr(24785)


Hi,
Let x represent the days it takes together
Amount done per day is the Equalizer
1/4 + 1/6 + 1/2 = 1/x
Mulitplying both sides of the equation by 12x so as all denominators = 1
3x + 2x + 6x = 12
11x = 12
x = 12/11
x = 1.09 hrs